OpenStudy (anonymous):

100 miles in 2.5 hours in unit rate ????

OpenStudy (phi):

they are asking how far you can go in 1 hour (unit means one here)

OpenStudy (anonymous):

yes

OpenStudy (phi):

you could write it as a ratio 100/2.5 = x/1 which says: 100 miles is to 2.5 hours as x miles is to 1 hour x/1 (x is the unknown rate) means x divided by 1, and is just x x = 100/2.5 simplify 100/2.5